The professor tells you on the first day of class that there will be five tests during the term. From the scores on these tests for each student, the professor will compute a measure of central tendency that will serve as the student’s final course grade. Before taking the first test, you the student must choose whether you want your final grade to be the mean or the median of the five test scores. Which would you choose? Why?
Answer #1

Please note that

Mean : mean is the sum of all observations devided by the number of observations

Median: median is the middle observation when the observations are arranged in increasing/decreasing order.

Suppose Jon get the marks in 5 tests ( out of 50) are 40, 42, 46, 47 and 39

So we get mean = (40+42+46+47+39) / 5= 214/5 = 42.8

So we get mean = 42.8

And for finding the median let's first arrange the data in increasing order. So we get arranged data-

39, 40, 42,46,47

So we get median = 42

So here we get mean > median

That is mean is greater than median.

Hence we choose Mean.

That is final grades to be the Mean.